I need to form the following equations and plot the graph of these equations
linear equation with one variable
Linear equation wit two variables.
: I need to form the following equations and plot the graph of these equations
linear equation with one variable
Linear equation wit two variables.

Answer by mangopeeler07(445)

Your variables will be x and y, since x and y are the variables in the coordinate plane. The equation for a line is y=mx+b, so just plug in values for m and b, and you will have the equation of a line with two variables. For the equation of a line with one variable, simply use x or y and assign a value. For example:

y=mx+b
Plug in any old thing for m and b:
y=3x+5

Now plot it:
Solved by pluggable solver: DESCRIBE a linear EQUATION: slope, intercepts, etc
Equation -3 x + 1 y = 5 describes a sloping line. For any
equation ax+by+c = 0, slope is -a/b = --3/1.
  • X intercept is found by setting y to 0: ax+by=c becomes ax=c. that means that x = c/a. 5/-3 = -1.66666666666667.
  • Y intercept is found by setting x to 0: the equation becomes by=c, and therefore y = c/b. Y intercept is 5/1 = 5.
  • Slope is --3/1 = 3.
  • Equation in slope-intercept form: y=3*x+5.
graph( 500, 500, -1.66666666666667-8, -1.66666666666667+8, 5-8, 5+8, 3*x+5 )



Now take x or y (we will do both for examples):
y
and assign a value
y=4

Now plot it:
Solved by pluggable solver: DESCRIBE a linear EQUATION: slope, intercepts, etc
This equation defines a horizontal line. There is no X intercept. Y intercept is 4/1=4
graph( 500, 500, -10, 10, 4 - 10, 4+10, 4/1 )



Now take x
x
and assign a value
x=2

Now plot it:
Solved by pluggable solver: DESCRIBE a linear EQUATION: slope, intercepts, etc
this equation defines a vertical line.

  • X intercept: 2. X intercept is the intersection of the line with the x axis.

  • Y intercept: not defined (the line IS vertical)

  • Slope: not defined, can be thought of as infinite

graph( 500, 500, 2-5, 2+5, -10, 10, 1000000x-1000000*2 )



That's all there is to it.